# Mitigating Perspective Distortion-induced Shape Ambiguity in Image Crops

### v1

@@ 1 Introduction
However, if we consider crops around the object, say those output by an object detector, and try making a prediction for the circle’s distance based on the appearance of the crop, this can prove to be quite difficult. Fig. 1 shows fixed sized center crops around the center of the circle’s image. It turns out, just looking at the crops while ignoring the location of the crop, it is no longer possible to predict the distance of the circle from the camera! A 0.5 cm wide image could correspond to Circle A that is 2 cm away directly in front of the camera, or Circle B that is further away (2.43 cm away) but off to the side. This simple example only exposed global pose error, but our detailed case study in Sec. 3 reveals that this ambiguity exists both in global pose and root-relative 3D shape. We call this Perspective Distortion-induced Shape Ambiguity in Image Crops or PSAC.
@@
(a.1) Root-relative 3D (a.2) Visualization of parallelepipeds with high root-relative 3D error and low centered 2D keypoint error.
Refer to caption Refer to caption
(b.1) Absolute 3D (b.2) Visualization of parallelepipeds with high absolute 3D error and low centered 2D keypoint error.
Refer to caption Refer to caption
Figure 2: Parallelepipeds visualization. Figure (a.1) plots the root relative 3D keypoint error vs. 2D keypoint error of different parallelepipeds placed at different locations in the camera’s field of view w.r.t. a reference cuboid shown in the green box in (a.2). Points are color coded with the distance between the parallelepipeds crops. As we let crops go farther away (the red points), we start finding parallelepipeds that have very different 3D shape but happen to project such that their 2D keypoints look the same as the reference cuboid. Figure (a.2) shows these 3D parallelepipeds in the top row and their renderings in the bottom row. The 1st figure in the green box is the reference w.r.t. which we measure 2D and 3D keypoint errors. Figures (b.1) and (b.2) present analogous plots where we measure absolute 3D error (rather than root-relative 3D error).
@@ 2 Related Work
Ambiguity in 3D from a single image: Scale-depth ambiguity is well known and is commonly dealt with by only evaluating 3D predictions up to a scale [29] or after a rigid alignment with the ground truth [21]. Where metric 3D predictions are needed, training is either done using domain specific knowledge [4], camera intrinsics as input [6] or information from predicted camera parameters [20, 23]. Notably, due to this ambiguity, models can not trivially be trained with scale and crop augmentations unless some adjustments are made [6, 9]. Ambiguity induced due to crops (the focus of our work) has not been analyzed in literature to the best of our knowledge. Note that this ambiguity is an artifact of planar perspective projection in images and doesn’t exist in spherical perspective projection. The differences and relationship between planar and spherical perspective have been used to simplify analysis, e.g. [26].
Geometric embeddings: Past works have considered the use of different encodings, based on pixel locations [25, 8], camera intrinsics [11] or extrinsics [37], as input to networks to improve performance. Liu et al. [25] show that preserving spatial location of the pixels helps in generative modelling, object detection, and reinforcement learning. Positional encodings, e.g. learned [8], relative [33], or rotational [32], are also commonly use in training transformer [35]-based models. Recent works [11, 15] have also studied the effectiveness of pixel-level encodings in multi-dataset setting (with varying cameras) on depth estimation tasks. Researchers have also considered canonicalizing the inputs (captured with different cameras) to a common camera intrinsics [6, 1] to indirectly provide camera information. [37, 14, 27] employ this idea in multi-view settings. Yifan et al. [37] used such encodings to inject geometric inductive biases (e.g. the epipolar geometry constraint) into a general perception model for multi-view stereo. Guizilini et al. [14] extend this further and use per-pixel embeddings to capture the image coordinate and camera intrinsics & extrinsics for multi-view depth estimation. Miyato et al. [27] propose a geometry-aware attention mechanism to encodes the geometric structure of tokens in transformers for novel view synthesis tasks.
Our use of the location of the crop in the camera’s field of view as an additional input to neural networks is motivated by these past works. Orthogonal to these works, our contribution is to demonstrate the effectiveness of this encoding in mitigating the perspective distortion-induced shape ambiguity in image crops.
@@ 3 Parallelepipeds Case Study
However, if the geometry of the object is also unknown, then it may no longer be possible to disentangle shape & pose from a given image crop. Consider the family of 3D shapes with a square face (of width $0.2$m lying in XY plane) that is extruded along an arbitrary direction, i.e. parallelepipeds with atleast one face being a square. Fig. 2 (a.2 top row) shows different parallelepipeds when placed directly in front of the camera (i.e. from one consistent viewing angle across the row). Note that the parallelepipeds here do not extend purely in the Z-direction (except for the leftmost reference in green box), but can extend at an oblique angle. Fig. 2 (a.2 bottom row) shows these same parallelepipeds when placed at different locations in the visual field. These locations were mined separately for each so as to maximize the visual similarity to the leftmost reference parallelepiped. As evident, even though the parallelepipeds across the row have different shapes, they all look the same in the 2D image (with a sub-pixel reprojection error), albeit for their 2D locations.
Figure 3: Predicting root relative (left) or absolute (right) 3D shape from 2D image crops fails in the absence of information about the location of the crop in camera’s field of view. Training loss saturates at a high value because of the inherent ambiguity. Adding in information about the location of the crop in camera’s field of view alleviates this ambiguity, leading the network to much better metrics for both root-relative and absolute 3D prediction.
This ambiguity only exists if we only look at the image crops locally without factoring in the absolute 2D location of the crop. Let’s treat the parallelepiped in the first column (a cuboid) as reference. Fig. 2 (a.1) plots the root relative 3D shape error as a function of the root relative 2D keypoint error of different parallelepipeds w.r.t. the reference cuboid. For root relative error, we compute the keypoint error after the centers of the parallelepipeds have been aligned. Points are color coded based on the 2D distance of the image crops from the cuboid crop (blue means image crops are closer, red means they are farther away). If we limit reasoning to only close-by crops, i.e. the blue points, low 2D keypoint error implies low 3D shape error. However, if we look at the farther away points (the red points), image with a low relative 2D keypoint error can actually have a high 3D shape error. This is the inherent ambiguity: even if two shapes appear similar visually, their 3D shapes may be vastly different. Fig. 2 (b.1) and (b.2) present analogous plots where we measure absolute 3D error instead of root-relative error.
@@ 4 Intrinsics-Aware Positional Encoding (KPE)
We mitigate the ambiguity induced by perspective distortion (Fig. 1) through design of encodings that capture the location of the crop in the camera’s field of view, referred to as Intrinsics-Aware Positional Encodings (KPE). Specifically, let’s denote the principal point for the original image by $(p_{x},p_{y})$ and focal length by $(f_{x},f_{y})$. For any pixel $(x,y)$, we can describe its position in the camera’s field of view via $\theta_{x}=\tan^{-1}\left(\frac{x-p_{x}}{f_{x}}\right)$ and $\theta_{y}=\tan^{-1}\left(\frac{y-p_{y}}{f_{y}}\right)$. Our proposal is to retain this information for pixels when images are cropped or resized for feeding into neural networks.
KPE can be used in different ways (Fig. 4). For a dense prediction task, e.g. depth estmation, we can concatenate this encoding in a dense manner to an intermediate feature map (or also to input pixels). For tasks that involve summarizing the geometric properties for objects, e.g. predicting its 3D bounding box or 3D pose, we consider a sparse variant, which encodes the center and corners of the relevant region of interest. We use both these variants in our experiments.
@@ 5 Experiments
We consider 3 challenging settings, involving the prediction of 3D geometric attributes from a single image, to show the effectiveness of KPE: 3D pose estimation of articulated objects in contact, dense depth prediction and 3D object detection. These tasks constitute both sparse and dense prediction of 3D attributes. Since our goal is to show the utility of KPE (not to achieve state-of-the-art results) we select some representative models from recent works on each of these tasks and incorporate KPE into their architectures.
Method Object Contact Motion
[…]
val split ArcticNet-SF [12] 8.0 59.0 44.1 36.8 11.8 11.3
+ KPE sparse 5.9 71.5 39.4 29.7 9.3 8.7
test split ArcticNet-SF [12] 6.4 53.9 44.7 36.2 11.8 9.1
+ KPE sparse 5.2 64.1 36.5 28.3 10.5 7.6
Table 1: 3D pose estimation of articulated objects in contact. We compare with ArcticNet-SF on multiple geometric tasks: 3D pose, contact and 4D motion estimation for articulated objects on ARCTIC. Our intrinsics-aware positional encoding leads to significant improvements across all metrics (See Sec. 5.1 for details). We use the sparse variant of KPE for these experiments.
5.1 3D Pose of Articulated Objects in Contact
Task: Given an egocentric image showing an articulated object in interaction, the goal is to estimate the 3D pose of the object in contact. This setting often involves occlusions induced by hands which makes it quite challenging. In such cases, additional cues about the spatial location or perspective distortion may provide a useful signal for pose prediction. Specifically, the 3D pose is represented using rotation in the camera coordinate frame, the translation of the root of the object with respect to the camera center and the angle of articulation between the bottom and top components of the articulated object. The object mesh is assumed to be available. We also consider contact estimation and 4D motion reconstruction to show the versatility of KPE.
Method: We adopt the single-image ArcticNet-SF [12] architecture released with the ARCTIC benchmark [12]. The network takes a single image as input, which is processed by a ResNet50 [17] backbone to get a feature map of resolution $7\times 7\times 2048$. These features are then average pooled and passed to a decoder to predict the 7-dimensional pose, consisting of rotation, translation & articulated angle of the object. The decoder iteratively predicts these entities using a HMR [22]-style architecture. For iterative prediction, the rotation and translation are initialized as a 0-vector whereas the intial articulation angle is predicted using the global features from the convolutional backbone. The network outputs rotation in axis-angle format and uses the weak perspective camera model to estimate the translation [5, 22, 24, 38]. The entire network is trained end-to-end using an L2 loss between the predictions and the ground truth for 7-dimensional pose, 3D keypoints and 2D projection of 3D keypoints.
@@
Dataset: ARCTIC [12] is a recent dataset consisting of hands interacting with articulated objects in a free-form manner. We select this dataset since it has a large range of dexterous interactions compared to existing datasets [34, 16]. There are 2 settings in the dataset: 1.5M frames from 8 allocentric views & 200K frames from 1 egocentric view. We show results in egocentric setting since the perspective distortion is more prominent due to objects being closer to the camera. The released dataset does not contain ground truth for test set, which is evaluated separately on the online leaderboard. We compare with ArcticNet-SF in this setting.
Metrics: We follow the evaluation protocol of the ARCTIC benchmark [12] and report the following metrics:
  • AAE: Average Absolute Error between the ground truth degree of articulation and the prediction.
  • Success: percentage of predicted object vertices with L2 error $<5\%$ of the diameter of the object.
  • CDev: Contact Deviation is average distance between the predicted in-contact hand-object vertices and ground truth
  • MRRPE: Mean Relative-Root Position Error between the predicted root keypoint location and the ground truth.
  • MDev: Between consecutive frames, Motion Deviation measures the disagreement in the direction of the movement of in-contact hand-object vertices.
  • ACC: Acceleration error is the difference in the predicted acceleration of the vertices and the ground truth.
Results: Adding KPE to ArcticNet-SF leads to consistent improvements across the 3D pose, contact and 4D motion estimation metrics (Tab. 1). We also ablate the sparse and dense variants of our encoding in Tab. 4. The dense variant is worse than the sparse variant, which could be due to the task requiring sparse predictions and the network getting confused by the spatial information about distractors around the object. In terms of design choices, we find that inserting KPE at the $7\times 7$ feature map level works better than adding it at the input or at the global feature level.

### v2

# https://bit.ly/AmbiguityEnc

@@ Abstract
Keywords:
3D from single image perspective distortion shape ambiguity
@@ 1 Introduction
However, if we consider crops around the object, say those output by an object detector, and try making a prediction for the circle’s distance based on the appearance of the crop, this can prove to be quite difficult. Fig. 1 shows
@@
fixed sized center crops around the center of the circle’s image. It turns out, just looking at the crops while ignoring the location of the crop, it is no longer possible to predict the distance of the circle from the camera! A 0.5 cm wide image could correspond to Circle A that is 2 cm away directly in front of the camera, or Circle B that is further away (2.43 cm away) but off to the side. This simple example only exposed global pose error, but our detailed case study in Sec. 3 reveals that this ambiguity exists both in global pose and root-relative 3D shape. We call this Perspective Distortion-induced Shape Ambiguity in Image Crops or PSAC.
@@
Refer to caption Refer to caption Refer to caption
(a) Root-relative 3D (b) Absolute 3D (c) Example Ambiguous Parallelepipeds.
Figure 2: Parallelepipeds Case Study. Figure (a) plots the root relative 3D keypoint error vs. 2D keypoint error of different parallelepipeds placed at different locations in the camera’s field of view w.r.t. a reference cuboid shown in the green box in (a.2). Points are color coded with the distance between the parallelepipeds crops. As we let crops go farther away (the red points), we start finding parallelepipeds that have very different 3D shape but happen to project such that their 2D keypoints look the same as the reference cuboid. Figure (b) shows a similar plot but for absolute 3D keypoint error. Figure (c) shows these ambiguous 3D parallelepipeds in the top row and their renderings in the bottom row. The 1st figure in the green box is the reference w.r.t. which we measure 2D and 3D keypoint errors.
@@ 2 Related Work
Ambiguity in 3D from a single image: Scale-depth ambiguity is well known and is commonly dealt with by only evaluating 3D predictions up to a scale [29] or after a rigid alignment with the ground truth [21]. Where metric 3D predictions are needed, training is either done using domain specific knowledge [4], camera intrinsics as input [6] or predicted camera calibration parameters [20, 22], e.g. focal length [22] using appearance cues [20]. Notably, due to this ambiguity, models can not trivially be trained with scale and crop augmentations unless some adjustments are made [6, 9]. Ambiguity induced due to crops (the focus of our work) has not been analyzed in literature to the best of our knowledge. Note that this ambiguity is an artifact of planar perspective projection in images and doesn’t exist in spherical perspective projection. The relationship between planar and spherical perspective have been used to simplify analysis, e.g. [26].
Geometric embeddings: Past works have considered the use of different encodings, based on pixel locations [25, 8], camera intrinsics [11] or extrinsics [40, 38], as input to networks to improve performance. Liu et al. [25] show that preserving spatial location of the pixels helps in generative modelling, object detection, and reinforcement learning. Different forms of positional encodings, e.g. learnable [8, 23], cartesian spatial grid [19], sinusoidal [35, 27], implicit [37, 31], help incorporate spatial inductive biases in GANs [37]. Positional encodings, e.g. learned [8], relative [34], or rotational [33], are also commonly use in training transformer [35]-based models. Recent works [11, 14] have also studied the effectiveness of pixel-level encodings in multi-dataset setting (with varying cameras) on depth estimation tasks. Researchers have also considered canonicalizing the inputs (captured with different cameras) to a common camera intrinsics [6, 1] to indirectly provide camera information. [38, 15, 28] employ this idea in multi-view settings. Yifan et al. [38] used such encodings to inject geometric inductive biases (e.g. the epipolar geometry constraint) into a general perception model for multi-view stereo. Guizilini et al. [15] extend this further and use per-pixel embeddings to capture the image coordinate and camera intrinsics & extrinsics for multi-view depth estimation. Miyato et al. [28] propose a geometry-aware attention mechanism for transformers for novel view synthesis tasks.
Our use of the location of the crop in the camera’s field of view as an additional input to neural networks is motivated by these past works. Orthogonal to these works, our contribution is to demonstrate the effectiveness of this encoding in mitigating the perspective distortion-induced shape ambiguity in image crops. We conduct this study across 3 representative 3D from a single image tasks, and where appropriate, show comparisons to past embeddings: Cam-Conv [11] and sinusoidal image location embeddings [35].
@@ 3 Parallelepipeds Case Study
However, if the geometry of the object is also unknown, then it may no longer be possible to disentangle shape & pose from a given image crop. Consider the family of 3D shapes with a square face (of width $0.2$m lying in XY plane) that is extruded along an arbitrary direction, i.e. parallelepipeds with atleast one face being a square. Fig. 2 (c, top row) shows different parallelepipeds when placed directly in front of the camera (i.e. from one consistent viewing angle across the row). Note that the parallelepipeds here do not extend purely in the Z-direction (except for the leftmost reference in green box), but can extend at an oblique angle. Fig. 2 (c bottom row) shows these same parallelepipeds when placed at different locations in the visual field. These locations were mined separately for each so as to maximize the visual similarity to the leftmost reference parallelepiped. As evident, even though the parallelepipeds across the row have different shapes, they all look the same in the 2D image (with a sub-pixel reprojection error), albeit for their 2D locations.
Figure 3: Predicting root relative (left) or absolute (right) 3D shape from 2D image crops fails in the absence of information about the location of the crop in camera’s field of view. Training loss saturates at a high value because of the inherent ambiguity. Adding information about the location of the crop in camera’s field of view alleviates this ambiguity, leading to better metrics for both root-relative and absolute 3D prediction.
This ambiguity only exists if we only look at the image crops locally without factoring in the absolute 2D location of the crop. Let’s treat the parallelepiped in the first column (a cuboid) as reference. Fig. 2 (a) plots the root relative 3D shape error as a function of the root relative 2D keypoint error of different parallelepipeds w.r.t. the reference cuboid. For root relative error, we compute the keypoint error after the centers of the parallelepipeds have been aligned. Points are color coded based on the 2D distance of the image crops from the cuboid crop (blue means image crops are closer, red means they are farther away). If we limit reasoning to only close-by crops, i.e. the blue points, low 2D keypoint error implies low 3D shape error. However, if we look at the farther away points (the red points), image with a low relative 2D keypoint error can actually have a high 3D shape error. This is the inherent ambiguity: even if two shapes appear similar visually, their 3D shapes may be vastly different. Fig. 2 (b) presents the analogous plot where we measure absolute 3D error instead of root-relative error.
@@ 4 Intrinsics-Aware Positional Encoding (KPE)
We mitigate the ambiguity induced by perspective distortion (Fig. 1) through design of encodings that capture the location of the crop in the camera’s field of view, referred to as Intrinsics-Aware Positional Encodings (KPE). Specifically, let’s denote the principal point for the original image by $(p_{x},p_{y})$ and focal length by $(f_{x},f_{y})$. For any pixel $(x,y)$, we can describe its position in the camera’s field of view via $\theta_{x}=\tan^{-1}\left(\frac{x-p_{x}}{f_{x}}\right)$ and $\theta_{y}=\tan^{-1}\left(\frac{y-p_{y}}{f_{y}}\right)$, as shown in Figure 4 (a). Our proposal is to retain this information for pixels when images are cropped or resized for feeding into neural networks.
4.1 Using KPE
KPE is quite flexible and can be used in different ways. We specifically consider two variations: Dense KPE (Figure 4 b) and Sparse KPE (Figure 4 c).
Dense KPE: We compute $\theta_{x}$ and $\theta_{y}$ for each per-pixel in a dense manner. Rather than using the raw angles, we encode them via sinusoidal encodings (4 frequencies each) [35]. The spatial feature map can be input into the network along side the raw pixels. Note that dense KPE can be computed at different resolutions and can be concatenated with a spatial feature maps at different points in the network.
Sparse KPE: We compute $\theta_{x}$ and $\theta_{y}$ for the 4 corners and the center of the crop. As for dense KPE, we encode them via sinusoids. Resulting embeddings can be concatenated and fed into MLP layers (along with flattened features for crop) or can be concatenated with spatial feature maps via broadcasting.
We experiment with sparse and dense KPE for the 3 representative tasks and find that dense KPE works best for dense tasks, while sparse KPE works best for tasks that involve summarizing the geometric properties for objects, e.g. predicting its 3D bounding box or 3D pose.
@@ 5 Experiments
We consider 3 challenging settings, involving the prediction of 3D geometric attributes from a single image, to show the effectiveness of KPE: 3D pose estimation of articulated objects in contact (Section 5.1), dense depth prediction(Section 5.2) and 3D object detection (Section 5.3). Since our goal is to show the utility of KPE (not to achieve state-of-the-art results) we select some representative models from recent works on each of these tasks and incorporate KPE into their architectures. These tasks encompass sparse and dense prediction of 3D attributes, and the representative methods cover both convolutional and transformer architectures. Thus, these together comprehensively test KPE. In each of the following subsections, we describe the application (the representative method and relevant metrics), how we apply KPE, and the results we obtain. We summarize the key takeaways in Section 6.
5.1 Application 1: 3D Pose of Articulated Objects in Contact [12]
Task: Given an egocentric image showing an articulated object in interaction, the goal is to estimate the 3D pose of the object in contact. Specifically, the 3D pose is represented using rotation in the camera coordinate frame, the translation of the root of the object with respect to the camera center and the angle of articulation between the bottom and top components of the articulated object. The object mesh is assumed to be available. We also consider contact estimation and 4D motion reconstruction to show the versatility of KPE.
Method: We adopt the single-image ArcticNet-SF [12] architecture released with the ARCTIC benchmark [12]. The network takes a single image as input, which is processed by a ResNet50 [17] backbone to get a feature map of resolution $7\times 7\times 2048$. These features are then average pooled and passed to a HMR [21]-style decoder to predict the 7-dimensional pose, consisting of rotation, translation & articulated angle of the object. The network outputs rotation in axis-angle format and uses the weak perspective camera model to estimate the translation [5, 21, 24, 39]. The entire network is trained end-to-end using an L2 loss between the predictions and the ground truth for 7-dimensional pose, 3D keypoints and 2D projection of 3D keypoints.
@@ 5.1 Application 1: 3D Pose of Articulated Objects in Contact [12]
Dataset: ARCTIC [12] is a recent dataset consisting of hands interacting with articulated objects in a free-form manner. There are 2 settings in the dataset: 1.5M frames from 8 allocentric views & 200K frames from 1 egocentric view. We show results in egocentric setting since the perspective distortion is more prominent due to objects being closer to the camera. The released dataset does not contain ground truth for test set, which is evaluated separately on the online leaderboard. We compare with ArcticNet-SF in this setting.
Table 1: 3D pose estimation of articulated objects in contact – Main Result. We compare with ArcticNet-SF [12] on multiple geometric tasks: 3D pose, contact and 4D motion estimation for articulated objects on ARCTIC [12]. Our intrinsics-aware positional encoding leads to significant improvements across all metrics and compares favorably to other ways of encoding the location of the crop and past encodings [11].
Method Object Contact Motion
AAE Success CDevho MRRPEro MDevho Acc
() $\downarrow$ (%) $\uparrow$ ($mm$) $\downarrow$ ($mm$) $\downarrow$ ($mm$) $\downarrow$ ($m/s^{2}$) $\downarrow$
a) Validation Split (KPE helps w.r.t. other choices)
ArcticNet-SF (full image) [12] 8.0 59.0 44.1 36.8 11.8 11.3
ArcticNet-SF (crop) 7.0 66.0 74.3 63.0 19.3 10.9
Full Image with Binary Mask 6.8 66.8 75.6 60.9 17.8 10.8
Cam-Conv [11] 6.3 68.8 38.7 29.7 10.1 9.4
Image Location (dense) 6.1 69.2 40.4 31.6 10.1 9.2
Image Location (sparse) 6.2 67.9 39.1 30.0 10.1 9.3
KPE (sparse) 5.9 71.5 39.4 29.7 9.3 8.7
b) Test Split
ArcticNet-SF (full image) [12] 6.4 53.9 44.7 36.2 11.8 9.1
KPE (sparse) 5.2 -19% 64.1 +19% 36.5 -18% 28.3 -22% 10.5 -11% 7.6 -16.4%
Table 2: 3D pose estimation of articulated objects [12] – KPE Design Choices. See Sec. 5.1 for details.
Method Object Contact Motion
AAE Success CDevho MRRPEro MDevho Acc
() $\downarrow$ (%) $\uparrow$ ($mm$) $\downarrow$ ($mm$) $\downarrow$ ($mm$) $\downarrow$ ($m/s^{2}$) $\downarrow$
a) Validation Split (Sparse KPE design choices)
     KPE (sparse) 5.9 71.5 39.4 29.7 9.3 8.7
     KPE (sparse center only) 7.2 66.8 44.5 33.1 10.9 10.2
     KPE (sparse w/ glb feat) 6.2 69.8 37.2 30.6 9.7 9.2
b) Validation Split (Dense KPE design choices)
     KPE (dense) 6.2 71.4 37.7 29.3 9.9 9.4
     KPE (dense w/ input) 7.3 68.1 48.7 38.3 13.6 10.2
c) Validation Split (Sparse vs. Dense)
     KPE (sparse) 5.9 71.5 39.4 29.7 9.3 8.7
     KPE (dense) 6.2 71.4 37.7 29.3 9.9 9.4
Metrics: We follow the evaluation protocol of the ARCTIC benchmark [12] and report the following metrics: a) AAE: Average Absolute Error between the ground truth degree of articulation and the prediction; b) Success: percentage of predicted object vertices with L2 error $<5\%$ of the diameter of the object; c) CDev: Contact Deviation is average distance between the predicted in-contact hand-object vertices and ground truth; d) MRRPE: Mean Relative-Root Position Error between the predicted root keypoint location and the ground truth; e) MDev: Between consecutive frames, Motion Deviation measures the disagreement in the direction of the movement of in-contact hand-object vertices; f) ACC: Acceleration error is the difference in the predicted acceleration of the vertices and the ground truth.
Results: Adding KPE to ArcticNet-SF leads to consistent improvements across the 3D pose, contact and 4D motion estimation metrics across both validation (Tab. 1 a) and test sets (Tab. 1 b). Using full images as input, as done in [12], obtains worse performance due to low resolution on the object. Focusing on the object, either through masking or by inputting only a crop, helps. KPE allows focusing on the object while also retaining the location of the object in the visual field, leading to the best performance. KPE made 2 design choices: a) use of sinusoids to encode, and b) use of crop’s location in the camera field of view rather than the 2D image. Both aspects are important and KPE is better than Cam-Conv [11] that represents locations in camera’s field of view but without sinusoids, and also just using sinusoidal embeddings of the 2D location in image.
We also ablate design choices for where to inject sparse and dense KPE in Table 2 b and Table 2 b. It is beneficial to inject sparse KPE at earlier layers rather than at the very end with global features. We also find that encoding both the scale and location of the crop is better than just encoding the location. For dense KPE, inputting it alongside the input is worse than using it alongside the $7\times 7$ spatial feature map coming out from the ResNet50 backbne.
While both sparse and dense versions of KPE improve performance, the sparse one being slightly better (Table 2 c).
5.2 Application 2: Dense Metric Depth Prediction [32, 4]
Task: Given a single RGB image, the goal is to predict the distance of each pixel to the camera center along the optical axis.
Method: We adopt the ZoeDepth architecture [4]. It consists of a transformer-based MiDaS [29] backbone to estimate relative depth, which is passed to a domain specific decoder along with multi-scale image features from the backbone, to predict the metric depth for a specific dataset. The model uses a discretized representation of 3D space (into a volumetric grid, referred to as bins) to first output coarse depth (represented as the center of the bin), which is then refined by an MLP in subsequent layers (referred to as the MetricBins [4] module). The model is trained using the scale-invariant log loss (similar to [3]) with $384\times 512$ images without any cropping and scaling augmentations during training. There are two variants of the model: a) ZoeD-X-N, trained from scratch and b) ZoeD-M12-N, initialized from a MiDAS model pre-trained on $12$ datasets (also on $384\times 512$ images).
Table 3: Dense metric depth prediction on NYU (Main Result). KPE leads to improved depth estimation accuracy when trained and tested with cropping and scaling augmentations. KPE outperforms Cam-Conv [11], a past encoding specifically design for depth estimation. We see gains in both settings with and without pre-training of the MiDaS backbone.
Method REL $\downarrow$ RMSE $\downarrow$ $\log_{10}$downarrow
a) No Pre-training for MiDaS backbone (KPE helps and is better than other choices)
 ZoeD-X-N 0.093 0.360 0.042
 ZoeD-X-N + Cam-Conv [11] 0.098 0.382 0.042
 ZoeD-X-N + KPE (dense) 0.088 -5.4% 0.334 -7.2% 0.038 -9.5%
b) MiDaS backbone pre-trained on 12 datasets (KPE helps and is better than other choices)
 ZoeD-M12-N 0.088 0.337 0.040
 ZoeD-M12-N + Cam-Conv [11] 0.096 0.376 0.041
 ZoeD-M12-N + KPE (dense) 0.082 -6.8% 0.313 -7.1% 0.035 -12.5%
Table 4: Dense metric depth prediction on NYU (KPE Design Choices). We conduct these comparisons in the setting where the MiDaS backbone is pre-trained on 12 datasets. See Section 5.2 for more details.
Method REL $\downarrow$ RMSE $\downarrow$ $\log_{10}$downarrow
a) Dense design choices
 ZoeD-M12-N + KPE (dense, inside MiDaS backbone) 0.082 0.313 0.035
 ZoeD-M12-N + KPE (dense, after MiDaS backbone) 0.085 0.323 0.038
b) Sparse design choices
 ZoeD-M12-N + KPE (sparse, inside MiDaS backbone) 0.083 0.321 0.037
 ZoeD-M12-N + KPE (sparse, after MiDaS backbone) 0.087 0.331 0.038
c) Sparse vs. Dense
 ZoeD-M12-N + KPE (dense) 0.082 0.313 0.035
 ZoeD-M12-N + KPE (sparse) 0.083 0.321 0.037